OpenStudy (anonymous):

what is 13 • [18 ÷ (15 − 9)] equal

OpenStudy (anonymous):

\[13\times \frac{ 18 }{ 15-9 }\]

OpenStudy (anonymous):

so first i'd subtract so 15-9=6 then divide 18/6=? and then multiply

OpenStudy (anonymous):

thanks for help!

OpenStudy (anonymous):

your welcome...